# MLLM-Fabric: Multimodal Large Language Model-Driven Robotic Framework for Fabric Sorting and Selection

### v1

@@ Abstract
Choosing the right fabric is crucial to meet functional and quality requirements in robotic applications for textile manufacturing, apparel production, and smart retail. We present MLLM-Fabric, a robotic framework powered by multimodal large language models (MLLMs) for fabric sorting and selection. The system includes a robotic arm, a camera, a visuotactile sensor, and a pressure sensor. It employs supervised fine-tuning and multimodal explanation-guided knowledge distillation to accurately classify and rank fabric properties. To facilitate further research, we release a dataset of 220 unique fabric samples, including RGB images and synchronized visuotactile and pressure data. Experimental results show that our Fabric-Llama-90B model consistently outperforms pretrained vision-language baselines in both property ranking accuracy and selection reliability.
@@ I Introduction
Robotic fabric manipulation has seen growing interest in recent years, especially for tasks involving material identification, physical property estimation, and deformable object handling [1]. While visual and tactile sensing have improved perception capabilities, conventional classification methods often fail to meet real-world selection demands. For instance, fabrics labeled under the same category, such as “cotton,” may vary drastically in softness, elasticity, or thickness. These differences critically affect performance in downstream tasks such as folding, dressing, or tool interaction. Effective fabric selection requires reasoning over continuous material attributes and contextual task relevance. Visual-only methods struggle to capture compliance or friction-related cues, while existing multimodal approaches [2, 3, 4, 5] often rely on unimodal embeddings, lack physical property supervision, or treat fabrics as static object categories without supporting structured comparison or functional decision-making. These limitations underscore the need for new paradigms in robotic material understanding.
This work introduces a physically grounded and task-aware robotic framework for fabric selection, formulated as a property-specific pairwise comparison problem. Instead of assigning absolute class labels, the system learns to determine which of two materials better satisfies a given property, such as softness or elasticity. This relative reasoning supports nuanced, context-sensitive decisions that are critical in real-world applications like adaptive grasping or clothing recommendation, where functional behavior rather than class identity determines success. At the core of our system is a multimodal language model trained on synchronized RGB, tactile, and pressure data. It is supervised through pairwise preference learning to support physically grounded, interpretable comparison of material properties. We further apply explanation-guided distillation to enhance reasoning clarity and allow abstention in uncertain cases. To support this formulation, we construct a dataset of 220 real-world fabrics, each annotated with four physical attributes and paired with co-registered RGB images, GelSight sequential frames, and pressure values. This dataset enables reproducible benchmarking of property-level material comparison under realistic sensory conditions.
@@
  • A new problem formulation that recasts robotic material selection as a property-specific pairwise comparison task, enabling functional decision-making under ambiguity and task variability.
  • A multimodal large language model framework trained using pairwise preference supervision and explanation-guided knowledge distillation, which supports interpretable and abstention-aware predictions from tactile, visual, and force inputs.
  • A robotic system that performs property-level reasoning for deformable object selection in tasks where relative material behavior, such as softness or elasticity, is more informative than discrete class identity.
@@
A demonstration video is included in the submission. The code and full dataset will be made publicly available on HuggingFace and a project website upon publication.
@@ II Related Work
TABLE I: Comparison of representative methods across capabilities: fabric-specific focus (Fabric), multimodal fusion (MM), comparative reasoning (CR), ranking (RK), task awareness (TA), and use of multimodal language models (MLLM). Refers to the study titled Active Clothing
Method Fabric MM CR RK TA MLLM
@@ III Methodology
This section describes the methodology of the proposed fabric selection robotic system (Fig. 1), detailing the data acquisition, multimodal perception, model training, and inference processes that enable accurate fabric property prediction and selection.
@@ III-A Visual Data Processing and Fabric Localization
The overall system workflow is illustrated in Fig. 2. During system operation, an Intel RealSense D435 camera mounted on the robotic arm captures RGB images and depth data to provide the necessary visual information for fabric localization and segmentation. To facilitate fabric identification and ensure consistent data association, ArUco tags are attached to each fabric sample, serving as unique IDs for pose tracking on the hanger. This step uses standard computer vision techniques to identify and record each sample’s position and orientation in 3D space. Depth data along the Z-axis is integrated with ArUco-based tracking to enhance pose estimation and account for shape variations or hanger misalignments. This integrated process ensures reliable visual localization, supporting the subsequent multimodal perception pipeline.
Following localization, the Segment Anything Model (SAM) [9] performs pixel-level segmentation to isolate each fabric sample from the background and neighboring objects in the RGB image. The segmented images are resized to 224$\times$224 pixels and stored with structured metadata, comprising tuples $(F_{i},P_{i},I_{i})$, where $F_{i}$ denotes the fabric ID, $P_{i}\in\mathbb{R}^{3}$ represents the 3D position vector, and $I_{i}$ corresponds to the segmented RGB image. This structured data ensures consistency across sensing modalities and supports downstream multimodal reasoning and property prediction tasks.
@@ III-B Fabric Tactile Sensing and Force-labeled Tactile Data
A dual-sensor setup is used for fabric tactile sensing (Fig. 3A): the left fingertip of the Franka Hand mounts a GelSight Mini [10] for high-resolution visuotactile imaging, while the right fingertip holds a strain-based planar pressure sensor with a dome surface for direct force measurement. The layered structure appears in Fig. 3B.
Although the GelSight sensor can estimate force from surface deformation, its accuracy degrades significantly on fabrics due to sensitivity to local thickness, compliance [3], and device-level variation [10]. To avoid relying on such uncertain estimates, we use a strain-gauge pressure sensor (HZC-30B, Chengying Sensor, China) mounted beneath a rigid planar surface. The sensor has a rated range of 0–10 kg, sensitivity of 1.0–1.5 mV/V, nonlinearity within $\pm$0.3full scale (F.S.), and creep within $\pm$0.5F.S. It operates from $-30\,^{\circ}\mathrm{C}$ to $70\,^{\circ}\mathrm{C}$ and supports 200% F.S. overload.
The GelSight and pressure sensor are mounted on opposing gripper fingertips to compress the fabric simultaneously. By Newton’s third law, the pressure reading equals the normal contact force on the GelSight surface. This setup enables synchronized recording of force and deformation without relying on image-based inference. The resulting stream comprises force-labeled tactile data, with each GelSight frame aligned to a corresponding pressure value. We use “force” to refer to these pressure readings throughout. This alignment substitutes for GelSight force calibration and provides accurate ground-truth supervision during training.
@@ III-C Multimodal Data Collection and Fusion Strategy
We collect data from 220 fabrics using a Franka arm, gripper, RealSense D435, GelSight Mini, and a planar pressure sensor. Each fabric is pressed twice (10s each) with a moderate-speed approach and slow closure for stable contact. GelSight and pressure data are sampled at 25 Hz and time-synchronized (Fig.3C), pairing each frame with a force reading. Fig.3D shows RGB examples illustrating visual diversity. The dataset contains RGB images, GelSight sequences, and aligned pressure values. GelSight captures spatial deformation linked to texture and compliance but not force; the pressure sensor provides scalar readings. Fusion preserves both deformation context and contact intensity.
To integrate this heterogeneous information, we first extract feature embeddings from each modality: RGB and GelSight images are processed by a frozen visual encoder, while pressure signals are represented in a language-compatible form and embedded via the model’s input projection layers. The resulting embeddings are unified within a shared token space and jointly processed by a multimodal transformer. This early-fusion strategy allows the model to learn cross-modal cues during training and inference, consistent with practices in recent vision-language models that align diverse modalities through token-level reasoning [11].
Each fabric sample is also annotated by experts with categorical labels for elasticity, softness, thickness, and texture. These annotations support supervised fine-tuning and evaluation.
@@ III-E1 Post Hoc Reasoning
$Q_{\text{post hoc}}=\text{MLLM}({z},Q_{\text{inst}},\tilde{r})$ (4)
@@ IV Experimental Setup
This section describes the experimental configuration employed to assess the effectiveness of the proposed methods.
@@ IV-B Model Training and Evaluation Setup
Each fabric sample is annotated with a four-dimensional attribute vector $\mathbf{a}=[a_{1},a_{2},a_{3},a_{4}]\in\{0,1,2\}^{4}$, representing elasticity, softness, thickness, and surface texture. Attributes are discretized into three ordinal levels: 0 (low), 1 (moderate), and 2 (high). To enhance supervision clarity, we retain only samples labeled 0 or 2 for training, discarding moderate-level examples to avoid semantic ambiguity and improve contrast in pairwise learning.
Two evaluation settings are used. The seen-fabric setting selects 200 training fabrics to form 24,000 extreme-level comparison pairs. From these, 4,000 pairs are randomly sampled for training and 400 for validation, ensuring equal distribution across the four attributes. Fabric instances may recur, but pairwise combinations are not duplicated across sets. The unseen-fabric setting evaluates generalization using 20 additional fabrics entirely excluded from training. Pair construction follows the same protocol with strict fabric-level isolation to prevent overlap. To assess extrapolation beyond binary extremes, we further evaluate the model on fine-grained comparisons involving adjacent attribute levels, such as 0 versus 1 and 1 versus 2. These cases are excluded from training and used solely to test the model’s ability to reason over subtle differences.
To mitigate overcommitment in ambiguous comparisons [15], we incorporate an abstain option into all prompts. While this mechanism is designed to improve reliability under uncertainty, its impact is not the focus of our primary evaluation. As an initial exploration, we conduct a separate test on 50 curated duplicate fabric pairs excluded from training. Our fine-tuned models abstained in fewer than 50 percent of these cases, whereas GPT-4o abstained in over 90 percent, indicating more conservative decision behavior. These findings highlight the limitations of our current training setup and motivate future work on explicit duplicate-aware supervision.
We evaluate Llama3.2-Vision-90B, chosen for its strong multimodal reasoning, and DeepSeek Janus-Pro-7B, a lighter baseline. GPT-4o is excluded due to its proprietary nature and reliance on closed APIs, which hinder reproducibility and integration into real-time systems. Fine-tuning is conducted on high-performance hardware. Llama3.2-Vision-90B is trained on an NVIDIA A800 GPU (80 GB RAM) for 6.5 hours. Janus-Pro-7B and Llama3.2-Vision-11B are trained on an NVIDIA A6000 (48 GB RAM) for 3 and 5 hours, respectively.
@@ IV-C Evaluation Metrics
where $TP$, $TN$, $FP$, and $FN$ denote true/false positives and negatives, respectively. Mean Accuracy (mean ACC) is the mean ACC across the four fabric properties and reflects overall ranking performance. To quantify prediction balance, we define Prediction Skewness (SK) as:
@@ V Experimental Results
We evaluate MLLM-Fabric across modalities, learning strategies, and model configurations, and demonstrate its advantages in property prediction, generalization, and selection tasks compared to prior baselines.
@@ V-A Fabric Property Sorting Experiment
We analyze how data modalities, training strategies, and model scales affect the accuracy, confidence, and generalization of pairwise fabric property prediction.
@@ V-B Fabric Selection Experiment
Scenario Properties Fabric-Llama-90B (proposed) GPT-4o Llama-Vision-90B GPT-4o IT Llama-Vision-90B IT
Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection Sorting Acc Selection
Athletic wear
for lexibility
Elasticity (2)
Softness (2)
100% 33.30% 50% 50% 50%
Summer dresses
for children
Softness (2)
Elasticity (2)
83.30% 20% 0% 62.50% 62.50%
Light draping
fabric for
home decor
Thickness (1)
Softness (2)
66.70% 20% 50% 37.50% 37.50%
Durable fabric
for frequent-use
furniture
Texture (2)
Thickness (1)
62.50% 28.50% 62.50% 62.50% 62.50%
Protective fabric
for work
environments
Thickness (1)
Elasticity (0)
50% 50% 50% 25% 25%
Heavy-duty material
for robust
applications
Texture (2)
Thickness (2)
75% 33.30% 25% 37.50% 37.50%
Soft lining
fabric for
formal suits
Thickness (0)
Softness (2)
Texture (0)
80% 42.80% 30% 50% 50%
Lightweight smooth
fabric for
formal shirts
Thickness (0)
Softness (2)
Texture (0)
90% 45% 9.10% 75% 75%
Ultra-soft
lightweight fabric
for accessories
Softness (2)
Thickness (0)
50% 50% 16.70% 75% 75%
Thick, luxurious
fabric for
premium interiors
Thickness (2)
Elasticity (0)
Texture (2)
61.50% 23% 15.30% 41.60% 41.60%
TABLE IV: Fabric selection performance in ten real-world scenarios. The first three columns show MLLMs (Fabric-Llama-90B, GPT-4o, Llama-Vision-90B) operating on raw multimodal inputs. The last two columns represent interaction trajectory (IT) pipelines, where a CNN-LSTM baseline predicts symbolic attribute levels forwarded to a frozen MLLM. Attribute levels are encoded as 0 (low), 1 (moderate), and 2 (high). marks correct selections; indicates failures.
V-C CNN-LSTM Tactile Baseline with Interaction Trajectories
We implemented a tactile-only CNN-LSTM model [3] to classify fabrics into three levels (low, moderate, high) across four attributes. On a held-out set of 20 unseen fabrics, it achieved a mean accuracy of 58%, with lower performance on texture and softness. To evaluate symbolic reasoning, we constructed an interaction trajectory (IT) pipeline by pairing the CNN-LSTM with a non-finetuned MLLM (GPT-4o or Llama3.2-Vision-90B). Each model received only the predicted attribute levels and issued a selection decision for each scenario. As shown in Table IV, both IT pipelines solved 5 out of 10 scenarios (50% accuracy), and their predictions were highly consistent. This consistency suggests that selection outcomes were largely determined by the upstream classifier. The modular structure restricts access to raw sensory input and prevents joint reasoning over perception and decision stages. In contrast, our end-to-end MLLM framework achieved 80% accuracy, highlighting the benefit of integrated training with multimodal grounding.
These comparisons suggest that end-to-end multimodal learning is more effective than modular pipelines, especially when alignment between perception and reasoning is critical. They also highlight the limitations of using symbolic attribute prediction as a proxy, and motivate tightly coupled training for grounded selection tasks.

### v2

@@ Abstract
Choosing appropriate fabrics is critical for meeting functional and quality demands in robotic textile manufacturing, apparel production, and smart retail. We propose MLLM-Fabric, a robotic framework leveraging multimodal large language models (MLLMs) for fabric sorting and selection. Built on a multimodal robotic platform, the system is trained through supervised fine-tuning and explanation-guided distillation to rank fabric properties. We also release a dataset of 220 diverse fabrics, each with RGB images and synchronized visuotactile and pressure data. Experiments show that our Fabric-Llama-90B consistently outperforms pretrained vision-language baselines in both attribute ranking and selection reliability. Code and dataset are publicly available at https://github.com/limanwang/MLLM-Fabric.
@@ I Introduction
Robotic fabric manipulation has attracted increasing attention in tasks such as material identification, physical property estimation, and deformable object handling [1]. Although visual and tactile sensing have improved perceptual capabilities, conventional classification methods often fail to meet practical selection requirements. For instance, fabrics labeled as “cotton” may vary significantly in softness, elasticity, and thickness, which directly impact downstream tasks including folding, dressing, and tool interaction. Effective fabric selection demands reasoning over continuous material properties and context-specific relevance. Visual-only methods are limited in capturing compliance and friction cues, while prior multimodal approaches [2, 3, 4, 5] often depend on unimodal embeddings, lack supervision on physical attributes, or treat fabrics as static object categories, without enabling structured comparison or functional decision-making. These limitations highlight the need for a new paradigm in robotic material understanding.
As shown in Fig. 1, this work introduces a physically grounded and task-aware robotic framework that formulates fabric selection as a property-specific pairwise comparison problem. Rather than assigning absolute class labels, the system learns to determine which of two materials better satisfies a specified property, such as softness or elasticity. This relative reasoning supports context-sensitive and function-oriented decisions, which are essential for real-world applications like adaptive grasping and clothing recommendation, where material behavior is more critical than category identity. At the core of our system is a multimodal large language model trained on synchronized RGB, tactile, and pressure data. The model is optimized through pairwise preference supervision to enable interpretable, physically grounded material comparisons. Additionally, we apply explanation-guided knowledge distillation to enhance reasoning clarity and allow abstention in uncertain cases. To support this formulation, we construct a dataset of 220 real-world fabric samples, each annotated with four physical attributes and paired with co-registered RGB images, GelSight sequences, and pressure readings. This dataset enables reproducible benchmarking of property-level material comparison under realistic sensory conditions.
@@
  • A new problem formulation that recasts robotic material selection as a property-specific pairwise comparison task, supporting functional reasoning under ambiguity and task variability.
  • A multimodal large language model framework trained with pairwise preference supervision and explanation-guided knowledge distillation, enabling interpretable and abstention-aware predictions from tactile, visual, and force modalities.
  • A robotic system that performs property-level reasoning for deformable object selection, where relative material behavior provides more informative guidance than discrete class identity.
@@ II Related Work
TABLE I: Comparison of representative methods across capabilities: fabric-specific focus (Fabric), multimodal fusion ($M$), comparative reasoning (CR), ranking (RK), task awareness (TA), and use of multimodal language models (MLLM). Refers to the study titled Active Clothing
Method Fabric $M$ CR RK TA MLLM
@@ III-A Visual Data Processing and Fabric Localization
The overall workflow is illustrated in Fig. 2. An Intel RealSense D435 camera mounted on the robotic arm captures RGB and depth data to enable fabric localization and segmentation. Each fabric sample is labeled with an ArUco tag, providing a unique ID for pose tracking on the hanger. Computer vision techniques are used to estimate the 3D position and orientation of each sample. Depth data along the Z-axis is fused with ArUco-based tracking to improve pose estimation under shape variation and hanger misalignment, ensuring robust visual localization for downstream multimodal processing. After localization, the Segment Anything Model (SAM) [9] performs pixel-level segmentation to isolate fabric samples from the background and neighboring objects. Each segmented image is resized to $\times224$ pixels and stored along with metadata tuples $(F_{i},P_{i},I_{i})$, where $F_{i}$ is the fabric ID, $P_{i}\in\mathbb{R}^{3}$ is the 3D position vector, and $I_{i}$ is the segmented RGB image. This structured representation maintains modality alignment and supports subsequent multimodal reasoning and property prediction.
@@ III-B Fabric Tactile Sensing and Force-labeled Tactile Data
A dual-sensor setup is used for fabric tactile sensing (Fig. 3A): the left fingertip of the Franka Hand mounts a GelSight Mini [10] for high-resolution visuotactile imaging, while the right fingertip holds a strain-based planar pressure sensor with a dome surface for direct force measurement (Fig. 3B). Although the GelSight sensor can estimate force from surface deformation, its accuracy degrades significantly on fabrics due to sensitivity to local thickness, compliance [3], and device-level variation [10]. To avoid such uncertainty, we use a strain-gauge pressure sensor (HZC-30B, Chengying Sensor, China) mounted beneath a rigid planar surface. It has a 0–10 kg range, 1.0–1.5 mV/V sensitivity, $\pm$0.3% F.S. nonlinearity, $\pm$0.5% F.S. creep, and operates from $30\,^{\circ}\mathrm{C}$ to $70\,^{\circ}\mathrm{C}$ with 200% F.S. overload tolerance. The two sensors are mounted on opposing fingertips to compress the fabric simultaneously. By Newton’s third law, the pressure reading equals the normal force on the GelSight surface. This setup enables synchronized recording of force and deformation without relying on image-based force estimation. The resulting stream forms force-labeled tactile data, where each GelSight frame is aligned with a corresponding pressure value. We refer to these pressure readings as “force” throughout. This alignment replaces the need for GelSight force calibration and provides accurate ground-truth supervision during training.
@@ III-C Multimodal Data Collection and Fusion Strategy
We collect data from 220 fabrics using a Franka arm, gripper, RealSense D435, GelSight Mini, and a planar pressure sensor. Each fabric is pressed twice for 10 seconds, with moderate-speed approach and slow closure to ensure stable contact. GelSight and pressure data are sampled at 25 Hz and time-synchronized (Fig. 3C), pairing each frame with a force reading. RGB examples in Fig. 3D illustrate visual diversity. The dataset includes RGB images, GelSight sequences, and aligned pressure values. GelSight captures spatial deformation related to texture and compliance, while the pressure sensor provides scalar force measurements. Fusion preserves both deformation context and contact intensity. To integrate this heterogeneous data, we extract feature embeddings from each modality. RGB and GelSight images are processed by a frozen visual encoder, and pressure signals are converted into language-compatible embeddings via the model’s input projection layers. The resulting embeddings are unified in a shared token space and jointly processed by a multimodal transformer. This early-fusion strategy enables learning of cross-modal cues during training and inference, following recent vision-language models that align diverse modalities through token-level reasoning [11]. Each fabric is also annotated with categorical labels for elasticity, softness, thickness, and texture by human experts. These annotations support supervised fine-tuning and evaluation.
@@ III-E1 Post Hoc Reasoning
[…]
@@ IV-B Model Training and Evaluation Setup
Each fabric is annotated with a four-dimensional attribute vector $\mathbf{a}=[a_{1},a_{2},a_{3},a_{4}]\in\{0,1,2\}^{4}$, representing elasticity, softness, thickness, and texture. Attributes are discretized into three levels: 0 (low), 1 (moderate), and 2 (high). To improve supervision clarity, we retain only level-0 and level-2 samples for training, discarding moderate-level cases to avoid ambiguity and enhance contrast in pairwise learning.
We use two evaluation settings. In the seen-fabric setting, 200 fabrics are used to generate 24,000 extreme-level comparison pairs. From these, 4,000 pairs are sampled for training and 400 for validation, evenly distributed across attributes. Fabric samples may recur, but pair combinations are unique across sets. The unseen-fabric setting uses 20 held-out fabrics with the same protocol and ensures no overlap at the fabric level. To evaluate generalization beyond binary extremes, we also assess the model on fine-grained comparisons between adjacent levels (0 vs. 1 and 1 vs. 2), which are excluded from training. To handle uncertainty in ambiguous cases [15], we include an abstain option in all prompts. Though not a focus of primary evaluation, we test it on 50 curated duplicate pairs. Our fine-tuned models abstained in fewer than 50% of cases, while GPT-4o abstained in over 90%, indicating more conservative behavior. This suggests current training lacks duplicate-aware supervision and motivates future improvements.
We evaluate Llama3.2-Vision-90B for its strong multimodal reasoning, and DeepSeek Janus-Pro-7B as a lightweight baseline. GPT-4o is excluded due to its closed-source nature and reliance on non-reproducible APIs. Fine-tuning is performed on high-performance hardware: Llama3.2-Vision-90B is trained on an NVIDIA A800 (80 GB RAM) for 6.5 hours; Janus-Pro-7B and Llama3.2-Vision-11B are trained on an NVIDIA A6000 (48 GB RAM) for 3 and 5 hours, respectively.
@@ IV-C Evaluation Metrics
where $TP$, TNTN, FPFP, and $FN$ denote true/false positives and negatives, respectively. Mean Accuracy (mean ACC) is the mean ACC across the four fabric properties and reflects overall ranking performance. To quantify prediction balance, we define Prediction Skewness (SK) as:
@@ V Experimental Results
UniTouch-7B None 0 0 0 0 0
Octopi-13B None 23 0 2 0 6 34.5
Octopi-13B D-SFT 2 3 21 5 7.75 60.9
Octopi-13B MEKD 4 5 40 0 12.25 78.7
@@
Scenario & Properties Fabric-L90B GPT-4o LV90B CNN IT CNN IT* Octopi-13B
Sort Sel Sort Sel Sort Sel Sort Sel Sort Sel Sort Sel
[…]
Summer dresses for children (Softness 2, Elasticity 2) 83.3% 20% 0% 62.5% 62.5% 16.7%
[…]
Durable fabric for frequent-use furniture (Texture 2, Thickness 1) 62.5% 28.5% 62.5% 62.5% 62.5% 37.5%
[…]
Heavy-duty material for robust applications (Texture 2, Thickness 2) 75% 33.3% 25% 37.5% 37.5% 25%
[…]
Lightweight smooth fabric for formal shirts (Thickness 0, Softness 2, Texture 0) 90% 45% 9.1% 75% 75% 36.3%
[…]
Thick, luxurious fabric for premium interiors (Thickness 2, Elasticity 0, Texture 2) 61.5% 23% 15.3% 41.6% 41.6% 38.4%
TABLE IV: Fabric selection performance across ten real-world scenarios. Sort = property-specific pairwise ranking accuracy; Sel = final selection correctness. Fabric-L90B = proposed Fabric-Llama-90B model, LV90B = Llama-Vision-90B, CNN IT / IT* = interaction pipelines combining a CNN-LSTM tactile predictor with GPT-4o (CNN IT) or Llama-90B (CNN IT*) for downstream selection. Attribute levels are encoded as 0 (low), 1 (moderate), and 2 (high). = correct selection; = failure.
@@ V-B Fabric Selection Experiment
V-C Comparative Analysis of Baseline Methods
V-C1 Symbolic Interaction Pipeline with Tactile Classification
We implemented a tactile classification pipeline following prior work [3], where a CNN-LSTM model was trained to predict fabric attributes across three discrete levels: low, medium, and high. During deployment, the predicted attribute levels were either used in rule-based selection logic or parsed into prompts for a large language model, forming two variants: CNN IT and CNN IT*. The former employed GPT-4o for downstream reasoning, while the latter utilized LLaMA3.2-Vision-90B. Although the tactile classifier achieved 58% average accuracy on attribute prediction for unseen fabrics, the end-to-end scenario-level selection accuracy reached only 50%, as shown in Table IV. This modular pipeline lacked access to raw sensory inputs during decision making, and could not support abstention behavior or adaptive reasoning under uncertainty. These limitations indicate the suboptimality of decoupled systems for property-sensitive material selection.
V-C2 Compared with Conventional Vision-Tactile MLLMs
UniTouch [7] aligns tactile signals from multiple sensors with vision–language embedding spaces, enabling unified representations, zero-shot generalization, and descriptive generation from single-sample inputs. Its inference interface, however, is oriented toward representation alignment and description rather than explicit property-level comparison. The model does not support pairwise input, scalar ranking, or task-conditioned selection. In our evaluation, queries consistently produced generic descriptive outputs that could not be deterministically mapped to comparative or selection decisions, yielding 0% accuracy (Table II).
Octopi [8] supports attribute-based prompting and can be queried in a pairwise format, but its inference is essentially categorical, with predictions made per sample and then interpreted for comparison. As shown in Table II, the pretrained Octopi-13B achieved a mean accuracy of only 6%, with meaningful predictions limited to elasticity while thickness, softness, and texture were nearly zero. When fine-tuned using either direct supervision or explanation-guided knowledge distillation, the mean accuracy improved slightly to 12.25%, driven mainly by texture, but the prediction skewness remained extremely high (78.7). This reflects severe output imbalance and unstable comparative reasoning. Consistent with these limitations, Octopi’s downstream selection success rate reached only 40% across ten real-world tasks (Table IV).
TVL [6] trains a touch–vision–language encoder on large-scale pseudo-labeled data, supporting open-vocabulary recognition and adjective tagging from single-sample inputs. While useful for category alignment, its outputs are free-form or tag-based rather than scalar and cannot support property-specific pairwise ranking or task-conditioned selection. Mapping these outputs into our evaluation protocol would require ad hoc thresholds and vocabulary choices, introducing non-determinism. Therefore, TVL is treated as related work but excluded from quantitative comparisons.
V-C3 Discussion
These visuo-tactile systems emphasize single-sample inference, either through categorical prediction (Octopi), representation alignment (UniTouch), or open-vocabulary description (TVL), and thus lack mechanisms for property-specific pairwise scalar reasoning. Octopi, although able to accept pairwise prompts, remains limited by categorical inference, showing low accuracy, high skewness, and weak downstream selection performance. UniTouch and TVL similarly focus on descriptive or representational objectives without ranking or selection. By contrast, our framework is trained for property-level pairwise decisions under scalar attributes and supports abstention when evidence is insufficient, which improves agreement with expert judgments and the reliability of downstream selection.